OpenStudy (anonymous):

Please helllppppp!? :D Which of the below systems of equations matches the following graph? x – 3y = 9 2x + 3y = 0 x + 3y = –9 2x + 3y = 0 x – 3y = –9 2x – 3y = 0 x + 3y = 9 2x – 3y = 0

OpenStudy (precal):

you sort of need the graphs?

OpenStudy (anonymous):

I know.. i have one

OpenStudy (anonymous):

Can you post a picture of the graph then I can try to help?

OpenStudy (anonymous):

yes. one sec! :)

OpenStudy (anonymous):

Just take some points on the graph and substitute them in the equation. Whatever equations satisfy the point are your equations

OpenStudy (anonymous):

satisfy the point in the sense which give 0 on substitution of x and y co-ordinates in the place of x and y

OpenStudy (anonymous):

I don't know how to get the graph on here :/

OpenStudy (anonymous):

oh wait! its right there

OpenStudy (anonymous):

they both intersect at (2,3) so if we substitute these equations in the above equations none satisfy so none of the equations match the graph

OpenStudy (anonymous):

hmm.... thats not one of the answers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well, I'm confused then..

jimthompson5910 (jim_thompson5910):

The point of intersection is (3,2) not (2,3)

OpenStudy (anonymous):

yeah sorry I'm wrong

OpenStudy (anonymous):

in that case the last set is your answer try substituting 2 and 3 in the place of x and y

OpenStudy (anonymous):

So, would this be the answer? x – 3y = –9 2x – 3y = 0

OpenStudy (anonymous):

no x+3y=9 (3+3*2=9) and 2x-3y=o (2*3-3*2=0)

OpenStudy (anonymous):

ahh. i was going to chose that one